Let X =


1 2


3 4


; E =


a


b





Find each of the following.


If the operation cannot be done : state undefined operation.


a) XE


b) EX


c) XT X where XT


stands for the transpose of X

1
Expert's answer
2022-03-05T08:01:06-0500

Solution


Given that


"X=\begin{bmatrix}
 1 & 2 \\
 3 & 4
\end{bmatrix}" , "E=\begin{bmatrix}
 a \\
b 
\end{bmatrix}", and "X^T=\begin{bmatrix}
 1 & 2 \\
 3 & 4
\end{bmatrix}^T=\begin{bmatrix}
 1 & 3 \\
 2 & 4
\end{bmatrix}"


Then


(a)


"XE=\begin{bmatrix}
 1 & 2 \\
 3 & 4
\end{bmatrix}\begin{bmatrix}
 a \\
b 
\end{bmatrix}"


"XE=\begin{bmatrix}
 1\times a+ 2\times b \\
 3\times a+ 4\times b
\end{bmatrix}"


"XE=\begin{bmatrix}
 2+2b \\
 3+4b
\end{bmatrix}"



(b)


For "EX=\begin{bmatrix}
 a \\
b 
\end{bmatrix}\begin{bmatrix}
 1 & 2 \\
 3 & 4
\end{bmatrix}"


We can see that number of columns in the first matrix "E" is just one and the number of rows in the second column "X" is two. Hence this multiplication is not possible.



(c)


For,


"X^TX=\begin{bmatrix}
 1 & 3 \\
 2 & 4
\end{bmatrix}\begin{bmatrix}
 1 & 2 \\
 3 & 4
\end{bmatrix}"


"X^TX=\begin{bmatrix}
 1+9 & 2+12 \\
 2+12 & 4+16
\end{bmatrix}"


"X^TX=\begin{bmatrix}
 10 & 14 \\
 14 & 20
\end{bmatrix}"